OpenStudy (anonymous):

Will medal Is this equation parallel? 3x-y=-9 2y-6x=-2

OpenStudy (zarkon):

\[3x-y=-9\] \[-y+3x=-9\] \[-2(-y+3x=-9)\] \[2y-6x=18\]

OpenStudy (anonymous):

what?

OpenStudy (snowsurf):

@mdjjh16 Let me ask you this. Do you what makes the equations parallel to each other.

OpenStudy (anonymous):

Yes, the slopes have to be the same

OpenStudy (anonymous):

I don't know how to do this one. The others I got but I'm stuck

OpenStudy (snowsurf):

Yes. These equations are not in slope intercept form. So make the equations like y = mx + b. Then look at the m (the slope) value and see what you get.

OpenStudy (anonymous):

see thats the problem, im messing up with the negatives and positives

OpenStudy (snowsurf):

Okay. First move the x on the right side for both equations. Then isolate y by diving the coefficient through the equation. Example 3x-y=-9 -y = -3x -9 divide by -1 y = 3x + 9 Now do that to the other equation. Then compare the slopes you find.

OpenStudy (anonymous):

2y-6x=-2 2y=6x-2 y=3x-1?

OpenStudy (anonymous):

so they are parellel

OpenStudy (snowsurf):

Yes they are. They have the same slope.

OpenStudy (anonymous):

ok thenks